OpenStudy (anonymous):

how do you divide x+5/x+3 / x+6/x^2+9x+18 please help and explain

OpenStudy (anonymous):

Which line is the 'big' divider.

OpenStudy (anonymous):

= x+6

OpenStudy (anonymous):

the middle /

OpenStudy (anonymous):

sorry correct answer should be x+5

OpenStudy (anonymous):

will show work next

OpenStudy (anonymous):

A/B/C/D=AD/BC

OpenStudy (anonymous):

\[(x+5)/(x+3) \div (x+6)/(x^2+9x+18)=\] since \[x^2+9x+18 = (x+3)(x+6)\] then we have\[(x+5)/(x+3)\times(x+3)(x+6)/(x+6) =\] cancel (x+3) and (x+6) x+5 left! hope it helps!